OpenStudy (gabe):

The area of a square garden is 128 m². How long is the diagonal?

jhonyy9 (jhonyy9):



jhonyy9 (jhonyy9):

than now you know the length of sides of the square so what is x=8sqrt2 and you know that for calculi the diagonal just you need using pythagira"s theorem do you know it ?

jhonyy9 (jhonyy9):

d^2 =x^2 +x^2 d^2 =2x^2 d = xsqrt2 can you calculi now the length of diagonal ?

jhonyy9 (jhonyy9):

x = 8sqrt2 so just substitute this value of x in place of x in this equation of d d=xsqrt2 so x=8sqrt2 what will get - ? try it please - i will help you - please collaborate